# Modeling the Hive Mind: Aligning Large Language Models with Online Communities

### v1

@@ 5.2.1 Automatic Evaluation
We train a classifier to determine the community from which a tweet originated. We achieve this by finetuning Llama-3 using demonstrations with the following template “Instruction: From these communities: Pro Eating Disorder, Keto & Diet, Body Image, Anti Eating Disorder, Healthy lifestyle & Weight Loss, and Weight Loss Drugs; which community does this Tweet belong to? {Tweet} Response: {Community name}”. We randomly sample 3,000 original tweets from each community’s corpus $D_{i}$ and construct a total of 18,000 demonstrations for finetuning. We train the classifier using 95% demonstrations and use the remaining 5% to test, leading to a test accuracy of 0.74. Using this model, we classify the finetuned LLM-generated tweets in $\mathbb{D}^{ft}=\{D^{ft}_{i}\}_{i=1}^{n}$ and RAG LLM-generated tweets $\mathbb{D}^{RAG}=\{D^{RAG}_{i}\}_{i=1}^{n}$, leading to an accuracy of 0.75 and 0.59, respectively. These results indicate that the classifier trained on original tweets accurately recognizes the tweets generated by the finetuned LLM. However, it performs poorly on the tweets generated by the RAG LLM, demonstrating that the finetuned LLMs better capture community-specific linguistic characteristics.
@@
We compute the semantic embeddings of $D_{i}$, $D^{ft}_{i}$, and $D^{RAG}_{i}$ using BERTweet Nguyen et al. (2020). We then measure the distance between these embeddings using the Fréchet Inception Distance (FID) Heusel et al. (2017). This metric provides a quantitative measure of the semantic distance between two text corpora. We implement it using the IBM comparing-corpora package Kour et al. (2022). $FID(D_{i},D^{ft}_{i})$ and $FID(D_{i},D^{RAG}_{i})$ for different communities are shown in Table 1. We see that $FID(D_{i},D^{ft}_{i})$ is much smaller than $FID(D_{i},D^{RAG}_{i})$, implying that the finetuned LLM outputs are more semantically similar responses to the original posts compared to the RAG LLM.
@@
Pro-ED 1.18 0.48
Body Image 1.42 0.37
Keto & Diet 1.05 0.51
Anti-ED 1.00 0.52
@@
1.19 0.54
Weight Loss Drugs 1.04 0.40
Table 1: Fréchet Inception Distances (FID) (1) between human-written tweets $D_{i}$ and RAG LLM generated tweets $D^{RAG}_{i}$, and (2) between human-written tweets $D_{i}$ and finetuned LLM-generated tweets $D^{ft}_{i}$. A smaller distance indicates more similarity.
@@
Figure 3: Emotional agreement (a) between human-written tweets and RAG LLM-generated tweets, and (b) between human-written tweets and finetuned LLM-generated tweets. The differences in affective alignment between pairs within each community are statistically significant at a 95% confidence level.
We analyze the emotions of tweets using SpanEmo Alhuzali and Ananiadou (2021). For each tweet, SpanEmo returns a vector of confidence scores of eleven emotions: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, and trust. For a community $C_{i}$, we sum the emotion confidence vectors of all tweets (i.e., $D_{i}$, $D^{ft}_{i}$, or $D^{RAG}_{i}$) and normalize them, resulting in an emotion distribution vector $\mathbf{e}_{i}$. We then compute the agreement between $\mathbf{e}^{ft}_{i}$ and $\mathbf{e}_{i}$, and between $\mathbf{e}^{RAG}_{i}$ and $\mathbf{e}_{i}$. The emotional agreement is measured as one minus the Jensen-Shannon distance between the two distribution vectors. As illustrated in Figure 3, for most communities, $D^{ft}_{i}$ more closely resembles the emotional tone of $D_{i}$ compared to $D^{RAG}_{i}$. This demonstrates that finetuning LLMs can effectively capture the authentic emotional tone of posts from communities.
We use Detoxify Hanu and Unitary team (2020) to measure toxicity in tweets Rajadesingan et al. (2020); Sheth et al. (2022). For a tweet, Detoxify returns a value between 0 and 1 indicating the level of toxicity. We only include tweets with toxicity levels equal to or greater than 0.05 for clarity and to reduce noise. Figure 4 shows the distributions of toxicity scores of human-written tweets $D_{i}$, RAG LLM-generated tweets $D^{RAG}_{i}$ and finetuned LLM-generated tweets $D^{ft}_{i}$. We observe that the toxicity distribution of $D^{ft}_{i}$ matches more closely to that of $D_{i}$ compared to $D^{RAG}_{i}$ for most communities, and tweets from the anti-ED community have the highest toxicity.
Figure 4: Toxicity distributions across different communities of human-written posts, RAG LLM generated posts, and finetuned LLM generated posts.
@@ 5.2.2 Human Evaluation
An annotator with expertise in EDs on social media was presented with 300 triplets, 50 from each community, where a triplet consists of a community name, a RAG LLM-generated tweet $d^{RAG}_{i,j}\in D^{RAG}_{i}$, and a finetuned LLM-generated tweet $d^{ft}_{i,k}\in D^{ft}_{i}$. Both tweets in a triplet are on the same topic and from the same community. For each triplet, the annotator was asked to decide which tweet was more aligned with the given community, by referring to the following characteristics: mis/use of ingroup language, references to themes in underlying distribution (e.g. the Body Image community often references nudity), use of capitalization, and coherence of message. In 225 out of 300 triplets, the annotator chose $d^{ft}_{i,j}$ as a better match, indicating the finetuned LLM is more aligned with the community.
@@
Online ED communities pose significant risks by promoting and normalizing harmful behaviors Lerman et al. (2023). Harm and toxicity are distinct in online discourse where toxicity detection algorithms may mistakenly flag explicit yet harmless language as toxic Sánchez et al. (2024). We come up with a dimension tailored to this ED domain where we assess harm by focusing on the underlying semantic content, as opposed to surface-level style. Our goal is for the finetuned LLM, $f^{\prime}_{i}$, to accurately capture the level of harm within these communities.
@@
We sampled 360 tweets across six communities, with 20 each from $D_{i}$, $D^{RAG}_{i}$, and $D^{ft}_{i}$. Two annotators with ED expertise labeled these tweets based on two tasks: (1) determine whether a tweet is harmful, and (2) classify harmful tweets into one of three fine-grained categories—body image objectification, relationship to food and exercise, or self-disclosure. Annotators achieved a Cohen’s Kappa score of 0.384 for identifying if harm was present and 0.519 for classifying fine-grained harm categories, indicating fair to moderate agreement.
A tweet was assigned to a harm category if both annotators agreed. Out of 360 tweets, 34 were classified into harm categories across $\mathbb{D}$, $\mathbb{D}^{RAG}$, and $\mathbb{D}^{ft}$. Figure 5 shows the distribution of these categories, demonstrating that finetuned LLMs better replicate the distribution of harm found in the community’s conversations.
@@ 6 Case Study: Screening Online Communities for Eating Disorders
In §$\S$5, we demonstrated that the finetuned LLM learns a more accurate representation of the community than the baseline method. This motivates us to apply psychometric instruments designed to evaluate an individual’s risk of EDs to online communities to help uncover unhealthy body and eating concerns within them.
@@
We prompt finetuned LLMs to respond to questions on the SWED screener. To account for randomness, for each item on the SWED questionnaire, the finetuned LLM generates 50 responses. The responses (Table 8 in Appendix D.5) are aggregated using a majority vote for each question. The results, Table 2, indicate that the Pro Eating Disorder community exhibits the highest levels of body image concerns, followed by the Keto & Diet community. Furthermore, both communities meet all three criteria signaling a high risk of ED pathology, whereas responses of the Anti Eating Disorder community are consistent with a low risk of ED.
These findings align with our empirical observations. The content shared by the Pro Eating Disorders community glorifies thinness and includes tips to promote disordered behaviors and body dysmorphia. Conversely, the Anti Eating Disorder community is critical of the diet culture and people who glorify EDs. The relatively high-risk score of the Keto & Diet community is a concerning indicator that this community may serve as a gateway to EDs. In contrast, the Body Image community, which mostly posts about body positivity, has a low risk of EDs, as does the Healthy Lifestyle & Weight Loss community. Although the latter focuses on weight loss, it appears to achieve this goal through healthy behaviors.
@@ 7 Conclusion
We demonstrate that aligning LLMs to online communities helps create high-fidelity digital proxies, which can be queried to reveal the implicit mindsets of these communities. When applied to online diet and fitness communities, the method uncovers communities with unhealthy body image and dieting beliefs that put their members at risk of EDs. This is important, as harmful communities that indoctrinate users into extreme ideologies Schmitz et al. (2022) or glorify EDs and self-harm Goldenberg et al. (2022) often evade moderation by using coded language that is opaque to outsiders or obfuscate harmful content via coded language and misspellings Chancellor et al. (2016d); Cobb (2017); Bickham et al. (2024). Our method can help online platforms bridge these obstacles to create safe and supportive online communities.
@@
The anonymized version of our dataset may contain implicit biases reflecting societal prejudices or unequal representation of demographic sub-groups. More specifically, ED symptoms have a history of being under-diagnosed in African American and Hispanic adolescents, in part due to stereotypical representation of ED being Caucasian adolescent girls Gordon et al. (2002). This historical bias could be inadvertently learned by our model, resulting in discriminatory behavior. In our future work, we hope to evaluate the model’s fairness across different user groups, allowing us to properly mitigate dataset biases.
@@
In our study, we expect the finetuned LLMs to replicate harmful narratives from online communities. This is done solely to demonstrate that, through our alignment framework, LLMs can accurately capture the nuanced language of these communities, including harmful content. Our objective is not to create models with malicious intent. We strongly advise that any future replication of this work be conducted with extreme caution.
@@
Diagnosing psychiatric illness at the community level comes with the risk of falsely diagnosing some community members. This could lead to unjust actions against users, such as unwarranted bans or removal of content. Additionally, approximating community behavior inherently excludes minority group members. Simultaneously, anorexia is one of the deadliest mental health disorders [^4] https://www.state.sc.us/dmh/anorexia/statistics.htm and participation in online pro-ED spaces heightens one’s disease risk Mento et al. (2021). By evaluating psychiatric illness on the community level, we can identify toxic communities, helping content moderation experts deploy proper interventions to promote healthy and safe online environments. We encourage the use of human moderators to review and validate the decisions made by our model, particularly in cases with low confidence scores.
@@ B.1 Demonstration Template for LLM finetuning
What would you tweet?
@@ C.1 Topics for Creating Synthetic Tweets
C.2 Prompt Template for Tweet Generation by RAG LLM

### v2

@@ 5.2.1 Automatic Evaluation
We train a classifier to determine the community from which a tweet originated by finetuning Llama-3 using demonstrations with the following template “Instruction: From these communities: Pro Eating Disorder, Keto & Diet, Body Image, Anti Eating Disorder, Healthy lifestyle & Weight Loss, and Weight Loss Drugs; which community does this Tweet belong to? {Tweet} Response: {Community name}”. We randomly sample 3,000 original tweets from each community’s corpus $D_{i}$ and construct a total of 18,000 demonstrations for finetuning. We train the classifier using 95% demonstrations and use the remaining 5% to test, with test accuracy of 0.74. We classify the finetuned LLM-generated tweets in $\mathbb{D}^{ft}=\{D^{ft}_{i}\}_{i=1}^{n}$ and LLM-Context-generated tweets $\mathbb{D}^{context}=\{D^{context}_{i}\}_{i=1}^{n}$, leading to an F1 accuracy score of 0.53 and 0.40, respectively. These results indicate that the classifier trained on original tweets accurately recognizes the tweets generated by the finetuned LLM. However, it performs poorly on the tweets generated by the LLM-Context, demonstrating that the finetuned LLMs better capture community-specific linguistic characteristics.
@@
We compute the semantic embeddings of $D_{i}$, $D^{ft}_{i}$, and $D^{context}_{i}$ using BERTweet Nguyen et al. (2020). We then measure the distance between these embeddings using the Fréchet Inception Distance (FID) Heusel et al. (2017). This metric provides a quantitative measure of the semantic distance between two text corpora. We implement it using the IBM comparing-corpora package Kour et al. (2022). $FID(D_{i},D^{ft}_{i})$ and $FID(D_{i},D^{context}_{i})$ for different communities are shown in Table 1. We see that $FID(D_{i},D^{ft}_{i})$ is much smaller than $FID(D_{i},D^{context}_{i})$ for almost all communities, implying that the finetuned LLM outputs are more semantically similar responses to the original posts compared to the LLM-Context.
@@
Pro-ED 1.16 0.82
Body Image 1.25 0.74
Keto & Diet 1.19 0.50
Anti-ED 0.84 0.42
@@
1.11 0.82
Weight Loss Drugs 0.90 1.99
Table 1: Fréchet Inception Distances (FID) (1) between human-written tweets $D_{i}$ and LLM-Context generated tweets $D^{context}_{i}$, and (2) between human-written tweets $D_{i}$ and finetuned LLM generated tweets $D^{ft}_{i}$. A smaller distance indicates more similarity.
@@
Figure 3: Emotional agreement (a) between human-written tweets and LLM-Context-generated tweets, and (b) between human-written tweets and finetuned LLM-generated tweets. The differences in the emotional alignment between pairs within each community are statistically significant at a 95% confidence level.
We analyze the emotions of tweets using Demux Chochlakis et al. (2023). For each tweet, Demux returns a vector of confidence scores of eleven emotions: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, and trust. For a community $C_{i}$, we sum the emotion confidence vectors of all tweets (i.e., the ones $D_{i}$, $D^{ft}_{i}$, or $D^{context}_{i}$) and normalize them, resulting in an emotion distribution vector $\mathbf{e}_{i}$. We then compute the agreement between $\mathbf{e}^{ft}_{i}$ and $\mathbf{e}_{i}$, and between $\mathbf{e}^{context}_{i}$ and $\mathbf{e}_{i}$. The emotional alignment is measured as one minus the Jensen-Shannon distance between the two distribution vectors He et al. (2024b). As illustrated in Figure 3, for most communities, $D^{ft}_{i}$ more closely resembles the emotional tone of $D_{i}$ compared to $D^{context}_{i}$. This demonstrates that finetuning LLMs can effectively capture the authentic emotional tone of posts from communities.
We use Detoxify Hanu and Unitary team (2020) to measure toxicity in tweets Rajadesingan et al. (2020); Sheth et al. (2022). For a tweet, Detoxify returns a value between 0 and 1 indicating the level of toxicity [^4] We only include tweets with toxicity levels equal to or greater than 0.05 for clarity and to reduce noise.. Figure 4 shows the distributions of toxicity scores of human-written tweets $D_{i}$, LLM-Context-generated tweets $D^{context}_{i}$ and finetuned LLM-generated tweets $D^{ft}_{i}$. We observe that the toxicity distribution of $D^{ft}_{i}$ matches more closely to that of $D_{i}$ compared to $D^{context}_{i}$ for most communities, and tweets from the anti-ED community have the highest toxicity.
Figure 4: Toxicity distributions across different communities of human-written tweets, LLM-Context-generated tweets, and finetuned LLM-generated tweets.
@@ 5.2.2 Human Evaluation
An annotator with expertise in EDs on social media was presented with 300 triplets, 50 from each community, where a triplet consists of a community name, a LLM-Context-generated tweet $d^{context}_{i,j}\in D^{context}_{i}$, and a finetuned LLM-generated tweet $d^{ft}_{i,k}\in D^{ft}_{i}$. Both tweets in a triplet are on the same topic and from the same community. For each triplet, the annotator was asked to decide which tweet was more aligned with the given community, by referring to the following characteristics: mis/use of ingroup language, references to themes in underlying distribution (e.g. the Body Image community often references nudity), use of capitalization, and coherence of message. In 248 of 300 triplets, the annotator chose $d^{ft}_{i,j}$ as a better match, showing better alignment of finetuned LLM with the community.
@@
Online ED communities pose significant risks by promoting and normalizing harmful behaviors Lerman et al. (2023a). Harm and toxicity are distinct in online discourse where toxicity detection algorithms may mistakenly flag explicit yet harmless language as toxic Sánchez et al. (2024). We come up with a dimension tailored to this ED domain where we assess harm by focusing on the underlying semantic content, as opposed to surface-level style. Our goal is for the finetuned LLM, $f^{\prime}_{i}$, to accurately capture the level of harm within these communities.
@@
We sampled 60 tweets from each community (360 for all six communities), with 20 each from $D_{i}$, $D^{context}_{i}$, and $D^{ft}_{i}$. Two annotators with ED expertise labeled these tweets based on two tasks: (1) determine whether a tweet is harmful, and (2) classify harmful tweets into one of three fine-grained categories—body image objectification, relationship to food and exercise, or self-disclosure. Annotators achieved a Cohen’s Kappa score of 0.453 for identifying if harm was present, and 0.617 for classifying fine-grained harm categories, indicating fair to moderate agreement (see more details in Limitations).
A tweet was assigned to a harm category if both annotators agreed. Out of 360 tweets, 41 were classified into harm categories across $\mathbb{D}$, $\mathbb{D}^{context}$, and $\mathbb{D}^{ft}$. Figure 5 shows the distribution of these categories, demonstrating that finetuned LLMs better replicate the distribution of harm found in the community’s conversations.
@@ 6 Case Study: Screening Online Communities for Eating Disorders
In §$\S$5, we demonstrate that the finetuned LLM learns a more accurate representation of the community than the baseline in-context learning method. This motivates us to apply psychometric instruments designed to evaluate an individual’s risk of EDs to online communities to help uncover unhealthy body and eating concerns within them.
@@
We prompt finetuned LLMs to respond to questions on the SWED screener. To account for randomness, for each item on the SWED questionnaire, the finetuned LLM generates 50 responses. The responses (Table 9 in Appendix D.5) are aggregated using a majority vote for each question. The results, as presented in Table 2, indicate that the Pro Eating Disorder community exhibits the highest levels of body image concerns, followed by the Keto & Diet community. Furthermore, both communities meet all three criteria signaling a high risk of ED pathology, whereas responses of the Anti Eating Disorder community are consistent with a low risk of ED.
These findings align with our empirical observations. The content shared by the Pro Eating Disorders community glorifies thinness and includes tips to promote disordered behaviors and body dysmorphia. Conversely, the Anti Eating Disorder community is critical of the diet culture and people who glorify EDs. The relatively high-risk score of the Keto & Diet community is a concerning indicator that this community may serve as a gateway to EDs. Experts caution that the keto diet’s emphasis on restrictive dieting eliminating carbs could expose vulnerable individuals to binge eating disorder and anorexia Kelvas (2023); Polivy and Herman (1985). In contrast, the Body Image community, which mostly posts about body positivity, has a low risk of EDs, as does the Healthy Lifestyle & Weight Loss community. Although the latter focuses on weight loss, it appears to achieve this goal through healthy behaviors.
@@ 7 Conclusion
We demonstrate that aligning LLMs to online communities helps create high-fidelity digital proxies, which can be queried to reveal the implicit mindsets of these communities. When applied to online diet and body image communities, the method uncovers communities with unhealthy body image and dieting beliefs that put their members at risk of eating disorders. This is important, as harmful communities that indoctrinate users into extreme ideologies Schmitz et al. (2022) or glorify eating disorders and self-harm Goldenberg et al. (2022) often evade moderation by using coded language that is opaque to outsiders or obfuscate harmful content via coded language and misspellings Chancellor et al. (2016d); Cobb (2017); Bickham et al. (2024). As social data are increasingly abundant, our method is generalizable to study various online communities in different fields and can assist online platforms in overcoming challenges to foster safe and supportive environments.
@@
The anonymized version of our dataset may contain implicit biases reflecting societal prejudices or unequal representation of demographic subgroups. More specifically, ED symptoms have a history of being under-diagnosed in African American and Hispanic adolescents, in part due to stereotypical representation of ED being Caucasian adolescent girls Gordon et al. (2002). This historical bias could be inadvertently learned by our model, resulting in discriminatory behavior. In our future work, we hope to evaluate the model’s fairness across different user groups, allowing us to properly mitigate dataset biases.
@@
Low Inter-Annotator Agreement for ED Harm Categorization.
The annotators in our harm categorization achieve a Cohen’s Kappa score from 0.384 to 0.519, which indicates fair to moderate agreement Landis (1977). Since no prior work has specifically focused on categorizing harmful ED content, we develop a harm taxonomy in collaboration with psychologists and clinicians specializing in eating disorders. This ongoing process has introduced uncertainty in defining some categories, leading to annotation discrepancies. Additionally, content discussing eating disorders on social media can be nuanced and implicit—while some posts may appear benign, they can subtly normalize harmful behaviors or contain triggering details for vulnerable users, further complicating annotation.
@@
In our study, we expect the finetuned LLMs to replicate harmful narratives from online communities. This is conducted solely to demonstrate that, through our alignment framework, LLMs can accurately capture the nuanced and authentic language of these communities, including harmful content. Our objective is not to create models with malicious intent; however, we acknowledge the potential scenario of vicious actors exploiting this framework to extract, amplify, and regenerate harmful information from social data. To mitigate this risk, we will release our code only upon eligible and transparent requests, ensuring that it is shared responsibly with researchers who have legitimate purposes and adhere to strict ethical and legal guidelines. We strongly advise that any future replication of this work be conducted with the utmost caution to prevent misuse and protect against the spread of harmful content.
Individual-Level Diagnosis
Existing computational frameworks that diagnose or predict mental illness based on individuals’ online activity and content raise significant ethical concerns due to the lack of user consent. These approaches often analyze personal data—even when publicly accessible—and infer sensitive individualized information like mental well-being without explicit permission from the users. By aggregating users’ data for community-level diagnosis, we can address these privacy concerns without infringing on individual autonomy, allowing for valuable insights to inform community-wide policy creation.
@@
Diagnosing psychiatric illness at the community level comes with the risk of falsely diagnosing some community members. This could lead to unjust actions against users, such as unwarranted bans or removal of content. Furthermore, psychological profiling of online communities sets a precedence for a slippery slope of increasingly intrusive monitoring and potentially creates a chilling effect on free speech in these spaces. Community members will anticipate and normalize heavy surveillance and thus self-censor or withdraw entirely from community discussions due to fear of revealing sensitive information that could lead to unintended consequences such as involuntary interventions. This would severely harm mental health online spaces by undermining the core values of trust and community. Additionally, approximating community behavior inherently excludes minority group members. Simultaneously, anorexia is one of the deadliest mental health disorders [^5] https://www.state.sc.us/dmh/anorexia/statistics.htm and participation in online pro-ED spaces heightens one’s disease risk Mento et al. (2021). By evaluating psychiatric illness on the community level, we can identify toxic communities, helping content moderation experts deploy proper interventions to promote healthy and safe online environments. We encourage the use of human moderators to review and validate the decisions made by our model, particularly in cases with low confidence scores.
@@
Acknowledgements
This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112290021 and Air Force Ofice of Scientific Research (AFOSR) under grant No. FA9550-23-1-0551. The authors are grateful to Ellen Fitzsimmons-Kraft for providing access to the eating disorders screener and to Aryan Karnati for aid with data analysis.
@@ A.1 Search Terms
Disordered eating behaviors exist along a spectrum between normal eating patterns and clinically diagnosable eating disorders Pereira and Alvarenga (2007). Previous studies have shown that restrictive diets, such as keto (“ketodiet”) or intermittent fasting (“intermittenfasting”), often intended for weight loss, are linked to a heightened risk of developing EDs Elran-Barak et al. (2015); Ganson et al. (2022); Cuccolo et al. (2022). While these behaviors may not meet clinical diagnostic criteria, they can act as a gateway, steering individuals toward more harmful online communities that promote disordered eating Lerman et al. (2023b). Additionally, we wanted to explore discussions from the opposite perspective (“bodypositivity”), where critics of diet culture and advocates for positive body image and healthy eating raise their voices.
@@ B.1 Demonstration Template for LLM finetuning
Share your thought with a tweet.
@@ C.1 Topics for Creating Synthetic Tweets
C.2 Prompt Template for Tweet Generation by LLM-Context
@@ C.3 Demonstration Template for Tweet Origin Classification
C.4 LLM Tweet Generation
Table 8 shows examples of LLM generated tweets. These examples across different communities and topics demonstrate that the finetuned LLM generates tweets more aligned with the assigned topic, effectively capturing the community’s consensus, norms, and linguistic patterns. In contrast, the LLM-Context-generated tweets show less specificity and coherence with the community’s established discourse. This highlights the finetuned model’s superior ability to reflect the language and cultural context of the target group.
Figure 8 shows the distribution of ROUGE-L scores between tweets in a community’s synthetic corpus $D^{ft}_{i}$ or $D^{context}_{i}$ and their most similar tweets within the corpus. Figure 9 shows the distribution of perplexity scores of tweets in a community’s synthetic corpus $D^{ft}_{i}$ or $D^{context}_{i}$. Figure 10 shows the distribution of ROUGE-L scores between tweets in a community’s synthetic corpus $D^{ft}_{i}$ or $D^{context}_{i}$ and their most similar tweets in the community’s original corpus $D_{i}$.
Figure 8: Distribution of ROUGE-L scores between tweets in a community’s synthetic corpus $D^{ft}_{i}$ or $D^{context}_{i}$ and their most similar tweets within the corpus.
Figure 9: Distribution of perplexity scores of tweets in a community’s synthetic corpus $D^{ft}_{i}$ or $D^{context}_{i}$.
Figure 10: Distribution of ROUGE-L scores between tweets in a community’s synthetic corpus $D^{ft}_{i}$ or $D^{context}_{i}$ and their most similar tweets in the community’s original corpus $D_{i}$.
Community Topic In-context Finetuned
Anti Eating Disorder body dysmorphia some of yall on edtwt need to mind fkg business??? edtwt is so bizarre 2 me; how are you making abt how ugly you think people are? dont you hate yourself? id worry abt that first,
Body Image extreme diet everyone has something about their body they’re not completely happy with. don’t focus on that! love yourself, love your body, love your life. im down to 15 stone 12!!! so so proud of myself, i know its been a long journey but i kept going and now im reaping the rewards. change happens over time and patience is the key.
Weight Loss Drugs wegovy what is the best diet for weight loss? too many diet rules doesn’t work! wegovy has helped many overcome obesity and drop excess pounds. recently, the hashtag has been trending as a side effect of taking wegovy. however, is this a side effect of taking the medication or simply due to rapid weight loss? full article
Eating Disorder anorexia when his celebrity crush is my thinspo and my celebrity crush is his thinspo im going to the doctors soon so im gon na have to lose some weight before i go
Table 8: LLM generated outputs across different communities and topics.